OpenStudy (anonymous):

Betty is now twice as old is Jim. in 6 more years the sum of their ages will be 6 times what Jims age was 3 years ago. what was Jim's age last year

OpenStudy (loser66):

still need help?

OpenStudy (irishboy123):

solve: B = 2J & B + 6 + J + 6 = 6 (J -3), and then ans = J -1